Solve:
(x/(x+1)) - (x/(x-1))=-16/63
I tried to multiply each side by the LCD 63(x+1)(x-1) but then everything cancels each other out and I end up with 0=-16. What am I doing wrong?

Answer by stanbon(75887)

(x/(x+1)) - (x/(x-1))=-16/63
Multiply thru -by 63(x+1)(x-1) to get:
63x(x-1) - 63x(x+1) = -16(x^2-1)
-126x = -16x^2+16
16x^2-126x-16=0
8x^2-63x-8=0
x=[63+-sqrt(63^2-4*8*-8)]/16
x=[63+-sqrt4225]/16
x=8 or x=-1/8
====================
Cheers,
Stan H.